Please help me solve this problem
m=x+3y/2 find y

Answer by ankor@dixie-net.com(22740)

Please help me solve this problem
m = x + 3y/2; find y
:
Multiply equation by 2 to get rid of the denominator, results:
2m = 2x + 3y
:
2m - 2x = 3y; subtracted 2x from both sides
or
3y = 2m - 2x
:
y = (2/3)m - (2/3)x; divided both sides by 3